Need help: Suppose P(C) =

Need help:
Suppose P(C) = .048, P(M ∩ C) = .044, and P(M u C) = .524. Find the indicated probability.
P(M')
A) .528
B) .520
C) .480
D) .472
Answer #1
Suppose P(C) = .048, P(M ∩ C) = .044, and P(M u C) = .524. Find the indicated probability. P(M') A) .528 B) .520 C) .480 D) .472 ----------------- P(M or C) = P(M) + P(C) - P(M and C) Substitute to solve for P(M) -------- 0.524 = P(M) + 0.045 - 0.44 P(M) = 0.524 + 0.44 - 0.045 P(M) = 0.919 ===================== Cheers, Stan H.
answered by: jAVOR